Emssj:

A store receives a shipment of VCRs and CD players. A shipment of 5 VCRs and 4 CD players cost $1950. A shipment of 3 VCRs and 6 CD players cost $2250. find the cost of a VCR and the cost of a CD player

snowflake0531:

We have to write out a system of equations 2 equations we have 5 VSRs and 4 CDs is 1950 and 3 VSRs and 6 CDs is 2250 we can write that VSR is x, CD is y 5x+4y=1950 3x+6y=2250

snowflake0531:

So with this, youcould use desmos and graph and find the intersecting coordinate, and (x,y) would be your answers

Emssj:

This is how I need to set it up with 2 equations for each

1 attachment
snowflake0531:

Yes and write those equations 5x+4y=1950 3x+6y=2250

Emssj:

Then I just find the answer for both right ?

snowflake0531:

Yep, just graph it, it's easier that way xd

Emssj:

Okay thanks fo you think you can help me with the second one

Emssj:

Under the first problem in the photo

snowflake0531:

Can you copy and paste the question, it'll be easier lol

Emssj:

Yess

Emssj:

A basketball team stopped at a fast food restaurant after a game. They divided into two groups. One group bought 5 chicken sandwiches and 7 hamburgers for a cost of $24.90. The second group spent $28.80 and bought 5 chicken sandwiches and 9 hamburgers. What is the cost of each chicken sandwich and each hamburger

snowflake0531:

Okay, so let's first put x is chicken sandwiches y is hamburgers 5 chicken sandwiches and 7 hamburgers for a cost of $24.90. 5x+7y=24.9 $28.80 and bought 5 chicken sandwiches and 9 hamburgers 5x+9y = 28.8

snowflake0531:

So with these two equations, you can graph them on desmos and find x adn y

Emssj:

Where it says set up that means I put chicken sandwhiches right ?

snowflake0531:

yea x = chicken sandwiches, y = hamburgers

Emssj:

For 5x+9y=28.80

Emssj:

I got x=25 and y= 45 does that sound right for the answer ?

Emssj:

I used Photomath and that’s what it said was the answer but not sure

snowflake0531:

Did you put both equations in-? and are you talking about the first problem or second

snowflake0531:

1 attachment
snowflake0531:

for the first

1 attachment
Emssj:

Ohh no I did not I only put the equation for chicken sandwiches

Emssj:

Do u know how I would put the answer do I just put 2.25, 1.95

snowflake0531:

well, x is something, it's representing something so you would say something costs 2.25 dollars, something costs 1.95 dollars etc

Emssj:

Ohh okay so I would put x= 2.25 y= 1.95

snowflake0531:

yea sure

Emssj:

Jason sells two sizes of drinks (small and large). One costs $1.00 and the other costs $1.50. He knows he sold a total of 230 drinks for a total of $285 . How many of each type of drink did he sell?

snowflake0531:

x is small y is large x+y = 230 drinks x+1.5y = 285 dollars

Emssj:

jack has 5$ and 10$ bills in his wallet . He has 25 bills totaling $230 . How many of each type of bills does he have ?

snowflake0531:

How about you first try to do it yourself xd

Emssj:

X =5$ bill Y=10$ bill X+y=25 5x+10y=230

snowflake0531:

Yes, that's coorrect sorry I didn't get the notification e.e

Emssj:

Macy has a total of 100 dimes and quarters. If the total value of the coins is $14.05, how many of each type of coin does she have?

Emssj:

X= dimes Y=quarters X+y=100 10x+.25=14.05

Emssj:

Is that correct ?

snowflake0531:

for the second equation, it owuld be 0.1x+ .25x =14.05

snowflake0531:

Since dime is not 10 dollars, it's a tenth xd